OpenStudy (drummmlers):

Subtract. y+2/4y + 16 minus 2/y^2 + 4y

jimthompson5910 (jim_thompson5910):

I'll get you started. First factor the denominators to get \[\Large \frac{y+2}{4y+16}-\frac{2}{y^2+4y}\] \[\Large \frac{y+2}{4(y+4)}-\frac{2}{y^2+4y}\] \[\Large \frac{y+2}{4(y+4)}-\frac{2}{y(y+4)}\] So the LCD is 4y(y+4). We must get each denominator equal to the LCD in order to combine them. That means you multiply top and bottom of the first fraction by y to get \[\Large \frac{y+2}{4(y+4)}-\frac{2}{y(y+4)}\] \[\Large \frac{y(y+2)}{4y(y+4)}-\frac{2}{y(y+4)}\] and then you multiply top and bottom of the second fraction by 4 to get \[\Large \frac{y(y+2)}{4y(y+4)}-\frac{2}{y(y+4)}\] \[\Large \frac{y(y+2)}{4y(y+4)}-\frac{4*2}{4y(y+4)}\] and I'll let you take over from here.